(a) If 913 J of heat are transferred into the system, and the system does 157 J of work, compute the change in the internal energy, and interpret the sign of your result.

                             (b) Meals-Ready-to-eat (MREs) are military meals that can be heated on a flameless heater via the reaction Mg(s) + 2 H2O(l) --- > Mg(OH)2(s) + 2 H2(g)

                 (i) Calculate ∆Hº298,rxn

                        (ii) Calculate the number of grams of Mg(s) needed for this reaction to release enough energy to increase the temperature of 75mL of water from 21 to 79 oC. Assume dwater = 1 g/mL. Note: heat given off is proportional to the moles of Mg(s).
Answer #1

(a) From first Law of thermodynamics ,

Q + W = dE

(where Q = Heat transferred to System, W = work done by system ; dE = change in internal energy ).

Q = 913 J ; W = -157 J (as per IUPAC convention work done by system is expressed as -ve sign )

So, dE ( change in internal energy) = 913 J - 157 J = +756 J

+ve sign of change in internal energy indicates that Internal energy of System have increased, part of heat transferred to system is absorbed as internal energy.

(b)We have following reaction ;

Mg(s) + 2 H2O(l) \rightarrow Mg(OH)2(s) + 2 H2(g)

i.   ∆r298 of reaction :

r298 = Products ΣHºmReactants ΣHºm

r298= {Hºm( Mg(OH)2(s)) + 2*Hºm(H2(g))} − {Hºm( Mg(s)) + 2*Hºm(H2O(l) )}

m is the standard molar enthalpy of a species.

The standard enthalpies of formation of elements in their reference states are zero, thus

m( Mg(s) = Hºm(H2(g) = 0

we have (various sources) : Hºm(H2O(l) = −285.83 KJ /mol

m( Mg(OH)2(s) = −924.7 KJ / mol

putting values,

  ∆r298= { −924.7 KJ / mol+ 0} − {0+ 2* −285.83 KJ /mol} = -353.04 KJ/mol

(-ve sign indicates exothermic reaction )

ii) Grams of Mg(s) needed for this reaction to release enough energy to increase the temperature of 75mL of water from 21 to 79 oC

Specific heat capacity of water(C) = 1 cal /g-deg. = 4.184 J / g-deg

Mass of water = 75 g

Amount of heat required, Q = m*dT * C = 75 g *58 deg* 4.184 J / g-deg = 18.2 KJ

Since reaction of 1 mol of Mg(s) releases = 353.04 KJ

Thus moles of Mg(s) needed = 18.2 KJ / 353.04 KJ =0.052 moles

( Molar mass of Mg = 24.3 g /mol )

or   Grams of Mg(s) needed =  0.052 moles * 24.3 g /mol = 1.26 g